3. Consider the following hypothetical supply and demand schedules for the newest smartphone at USF.

Supply Schedule Demand Schedule

Price Quantity Price Quantity

700 10,000 700 1,000

650 8,000 650 2,000

600 6,000 600 3 000

550 4,000 550 4,000

500 2,000 500 5,000

a. What is the equilibrium price and quantity? Briefly explain using the table.

b. At a price of 500, is there a shortage or a surplus of smartphones? What is the size of the shortage/surplus? Briefly explain using the table.

c. At a price of 650, what is the size of the shortage/surplus? Briefly explain using the table.
Answer #1

3.
a. The equilibrium price is 550 and quantity is 4,000 as equilibrium occurs at the point where supply equals demand. Using the table, we can see that quantity supplied = quantity demanded = 4,000 at price 550.

b. At a price of 500, quantity supplied = 2,000 and quantity demanded = 5,000. So, there is a shortage as quantity demanded exceeds quantity supplied. Shortage = quantity demanded - quantity supplied = 5,000 - 2,000 = 3,000

c. At a price of 650, quantity supplied = 8,000 and quantity demanded = 2,000. So, there is a surplus as quantity supplied exceeds quantity demanded. Shortage = quantity supplied - quantity demanded = 8,000 - 2,000 = 6,000